Verify that 0.200 mol of a nonvolatile solute in 125 g of benzene (C6H6) produces a solution


whose boiling point is 84.2 °C.

1
Expert's answer
2021-11-23T13:47:03-0500

Number of moles of solute = 0.200 mol

Weight of solvent (Benzene) = 125 g = 0.125 kg

Molality of solution "= \frac{0.20}{0.125} = 1.6 \;m"

Elevation of boiling point "= \u0394T_{bp} = K_{bp} m_{solute}"

"K_{bp}" = Molal elevation boiling point constant

"K_{bp} = 2.53 \; °C/m" for benzene

"ΔT_{bp} = 2.53 \times 1.6 = 4.048 \;°C"

Boiling point of solution = Boiling point of pure solvent + Elevation in boiling point

= 80.1 °C + 4.048 °C = 84.15 °C

Boiling point of solution = 84.15 °C